OpenStudy (anonymous):

What is the product of 5the square root of 3. • 4the square root of 15.? Simplify if possible.

OpenStudy (owlcoffee):

\[(5 \sqrt 3 )(4\sqrt15)\] We will make use of the property: \[\sqrt{A} \sqrt{B}=\sqrt{AB}\] Wich only means that the product of two radicals is equal to the radical involving the product of the two original radical numbers. But we have to express the "5" as a radical in order to make this worth. But "5" is the square root of 25, and "4" is the square root of 16, so we can replace it: \[\sqrt{25}\sqrt{3}\sqrt{16}\sqrt{15}\] So, now, applying the property: \[\sqrt{25.3}\sqrt{16.15}\] And we can apply it yet again: \[\sqrt{25.3.16.15}\] \[\sqrt{18000}\]

OpenStudy (anonymous):

sorry here are the answers 60 square root 5 20 square root 5 60 square root 15 20 square root 15 I think its the last one